The Rolling Creek Textile Mill produces cotton denim. The fixed monthly cost is $21,000, and the variable cost per yard of denim is $0.45. The mill sells a yard of denim for $1.30.
a. For a monthly volume of 18,000 yards of denim, determine the total cost, total revenue, and profit.
a. Determine the annual break-even volume for the Rolling Creek Textile Mill.

Answer by richwmiller(17219)

y=mx+b
m=1.30-.45=85 for profit
m=1.30 revenue
m=.45 cost
y=21000
x=18000


.85*x =21000 breakeven 24706 yards MONTHly! multiply by 12 for YEARly
y=.45*18000 +21000 for total cost 29100 for 18000
y=1.30*18000 total revenue = money received 23400
y=23400-29100=-5700 loss NOT profit for the 18000 yards